given the function h(r)=9r+4
h(3)=?
h(-4)=?
h(6.9)=?

Answer by rfadrogane(214)

given the function h(r)=9r+4
a. h(3)=?
b. h(-4)=?
c. h(6.9)=?
Sol'n:
a. h(3)= 9(3)+4 = 31 ---answer
b. h(-4)= 9(-4)+4 = -32 ----answer
c. h(6.9)= 9(6.9)+4 = 66.1 ---answer